Week of 190819

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes



  • local: Alberto (monitoring), Andrei (databases), Enrico (CERN storage), Jiri (ATLAS), Julia (WLCG), Maarten (ALICE + WLCG), Olga (CERN computing), Vladimir (LHCb)
  • remote: Christian (NDGF), Christoph (CMS), Dave (FNAL), Di (TRIUMF), Marcelo (CNAF), Mike (ASGC), Pepe (PIC)

Experiments round table:

  • ATLAS reports ( raw view) -
    • stopped brokering analysis jobs to SL6 queues, offline in 1 month
    • in one month stop brokering production jobs to SL6 queues, offline in two months
    • reprocessing at T1's continues

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • MC, user jobs and data restripping.
      • Massive staging at all T1
    • Issues:
      • RAL:
        • GGUS:142350; still issues accessing files on ECHO. Under investigation.
      • CNAF:
        • Outage. DT extended until Wed.
      • IN2P3
        • Staging problem

Sites / Services round table:

  • BNL:
  • CNAF: We have today a crew of technicians from a company hired to provide manpower needed bring the system back. Some of the conductors that bring power to the data center must be replaced. Tomorrow the new UPS system is going to be put in place. We hope to put the system up by Wednesday.
  • EGI:
  • FNAL: ESNet outage on the 16th for about 2 hours affected external connectivity.
  • IN2P3:
  • KISTI:
  • KIT:
  • NL-T1:
  • NRC-KI:
  • OSG:
  • PIC:
  • RAL:

  • CERN computing services: NTR
  • CERN storage services: NTR
  • CERN databases: NTR
  • Monitoring:
    • sites have until Aug 23 to open tickets about the July SAM A/R reports
  • MW Officer:
  • Networks: NTR
  • Security:


Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r16 - 2019-08-19 - MaartenLitmaath
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback